OpenStudy (anonymous):

Is my answers right? if not what do i need to do? For questions 3-4, factor the quadratic expression completely, and find the roots of the expression. 3. (3 points) 20x2 + 13x + 2 (5x+2)(4x+1) 4. (3 points) 49x2 - 14x - 3 (7x+1)(7x-3)

OpenStudy (anonymous):

the factorization is correct ,know solved the factored forms for the roots: (5x+2)(4x+1)=0=>x=-2/5 or x=-1/4 (7x+1)(7x-3)=0=>x=-1/7 or x=3/7

OpenStudy (anonymous):

THANK YOU